OpenStudy (anonymous):

a(3tx-2b)=c(dx-2) solve for x

OpenStudy (anonymous):

3atx - 2ab = cdx - 2c 3atx - cdx = 2ab - 2c x(3at - cd) = 2(ab - c) x = 2(ab - c)/(3at - cd)

OpenStudy (anonymous):

thanks!